Grade 11Discuss with colleagues and IITians

A balloon is rising up with a velocity of 29 m/s. A stone is dropped

from it and it reaches the ground in 10 sec. Height of the balloon when the stone was dropped from it is ?

Nithul Krishna
12 Years agoGrade 11

1 Answer

Apoorva
ApprovedApproved Tutor Answer12 Years ago
790m. OR 290m sol:When we drop a stone from a ballon it has same velocity as that of the balloon. SO,for the balloon, velocity (u)=29 m/s; time (t)=10seconds: acceleration due to gravity(g)=10 By eqn of motion... s=ut+1/2at^2; we get s=790m, but if we don`t havet to take effect of gravity;s=290m